A cyclist traveled at arate of 32mph to visit a nearby town. The cyclist averaged 10 mph on hte return trip. If the round trip took 6.3 hours, find the distance to the nearby town.
Answer by tvandenberg(45)

The trip is "x" miles each way.
Since time = distance / speed,
x / 32 + x / 10 = 6.3
Let's do this without a calculator. Chose a common denominator, 160.
5x / 160 + 16x / 160 = 6.3
21x / 160 = 6.3
21x / 16 = 63
x = 63 * 16 / 21 = 3 * 16 = 48
The town is 48 miles away.